OpenStudy (anonymous):

The center of the circle that passes through the points (-2,3), (0,-3), and (4,-3) is________ a) (2,2) b) there is not enough data to find the center c) (1,2) d) (2,1)

OpenStudy (anonymous):

first we can find the eq of circle passing through three points, then can check the centre

OpenStudy (anonymous):

how do I do that?

OpenStudy (anonymous):

general eq of the circle is \[x^2+y^2+2gx+2fy+c=0\]

OpenStudy (anonymous):

substitute the three points in this eq we get three eqs in three unknows find the unknows g,f and c

OpenStudy (anonymous):

centre is \[\sqrt{g^2+f^2-c}\]

OpenStudy (anonymous):

for(-2,3) 4+9-4g+6f+c=0

OpenStudy (anonymous):

for(0,-3) 9-6f+c=0

OpenStudy (anonymous):

for(4,-3) 16+9+8g-6f+c=0

OpenStudy (anonymous):

-4g+6f+c=-13 -6f+c=-9 8g-6f+c=-25

OpenStudy (anonymous):

subtracting 1 n 3 -12g+12f=12 -g+f=1

OpenStudy (anonymous):

sub 1 n 2 -4g+12f=-4 g-3f=1

OpenStudy (anonymous):

adding the resultant eqs -2f=2 f=-1 which gives g=-2 n c=-3

OpenStudy (anonymous):

center=(-g,-f)=(2,1)

OpenStudy (anonymous):

so D is the right option:)

OpenStudy (anonymous):

m sorry the formula i wrote earlier is to find the radius

OpenStudy (anonymous):

thank you for your help

OpenStudy (anonymous):

u got all the lenghy computations?

OpenStudy (anonymous):

lengthy*

OpenStudy (anonymous):

for the most part

OpenStudy (anonymous):

what do g and f represent though?

OpenStudy (anonymous):

if there is some confusion, u can ask:)

OpenStudy (anonymous):

these r constants

OpenStudy (anonymous):

the eq i wrote is the general eq of the circle

OpenStudy (anonymous):

finding out g,f and c, we can find the eq of the circle

OpenStudy (anonymous):

okay thank you

OpenStudy (anonymous):

welcome:)